# Strategizing with AI: Insights from a Beauty Contest Experiment

### v1

@@ Abstract
A Keynesian beauty contest is a wide class of games of guessing the most popular strategy among other players. In particular, guessing a fraction of a mean of numbers chosen by all players is a classic behavioral experiment designed to test iterative reasoning patterns among various groups of people. The previous literature reveals that the level of sophistication of the opponents is an important factor affecting the outcome of the game. Smarter decision makers choose strategies that are closer to theoretical Nash equilibrium and demonstrate faster convergence to equilibrium in iterated contests with information revelation. We replicate a series of classic experiments by running virtual experiments with modern large language models (LLMs) who play against various groups of virtual players. We test how advanced the LLMs’ behavior is compared to the behavior of human players. We show that LLMs typically take into account the opponents’ level of sophistication and adapt by changing the strategy. In various settings, most LLMs (with the exception of Llama) are more sophisticated and play lower numbers compared to human players. Our results suggest that LLMs (except Llama) are rather successful in identifying the underlying strategic environment and adopting the strategies to the changing set of parameters of the game in the same way that human players do. All LLMs still fail to play dominant strategies in a two-player game. Our results contribute to the discussion on the accuracy of modeling human economic agents by artificial intelligence.
Keywords: beauty contest, guess the number, dominant strategy, AI, LLM
@@ 1 Introduction
The recent appearance of large language models (LLMs) has resulted in numerous attempts to substitute humans with generative agents in various settings (see, for example, Park et al. (2023)). The motivation is simple: in many economic activities, the use of algorithms sooner or later will be cheaper and more productive than the use of labor. Still, at this point, it is not clear to what extent LLM can simulate the human’s behavior. Several papers investigated the differences between LLM decisions and the participants of economic experiments. In the recent manuscript, Horton (2023) replicates several classic experiments with LLM players and advocates the use of LLMs as models for ordinary economic agents. Our paper contributes to this strand of literature by studying the behavior of LLMs in the classic Guess the number game. This game belongs to a wider class of Keynesian beauty contest experiments and is of particular importance because it is designed to study level-$k$ reasoning, the ability of a player to make a sequence of conclusions as a function of their level of sophistication. To what extent LLM behaves like a human in making sequential conclusions? To answer this question, we replicate a series of well-known experiments with human participants by asking LLMs to play against the same groups of competitors.
@@
The rules of the Guess the number game are as follows. A group of $n$ players simultaneously and independently choose a number between 00 and $100$. Denote by $m$ the mean of all strategies played. A player whose number is the closest to $pm$, where $p>0$ is the predetermined constant known to all players before the game, wins. In case of a tie, all tied players get the corresponding share of the prize. For $p\in(0,1)$, there is a unique Nash equilibrium in the model where all players choose 00. In a particular case of $n=2$, choosing 00 is a weakly dominant strategy.
@@
In a particular case of $n=2$, one could possibly anticipate a significant share of players choosing 00, a weakly dominant strategy. However, this is not the case. Only $10\%$ of undergraduate students and $37\%$ of the audience of economics or psychology decision-making conferences chose 00 (Grosskopf & Nagel, 2008). Also, the mean of the numbers chosen by the professionals (22) is lower than the mean of the numbers chosen by the students (36). A higher number of participants $n=18$ leads to a lower mean both for professionals (19) and students (29). However, in the case of professionals, this difference is not statistically significant (Grosskopf & Nagel, 2008).
@@
Cognitive ability is also an important determinant of the outcome of the Guess the number game. Brocas & Carrillo (2020) designed a variant of the Guess the number game and demonstrated that the equilibrium behavior increases significantly between 5 and 10 years of age (from 17.9% to 61.4%) and stabilizes afterwards. Back to the claasic variant of the game, players with higher scores in a cognitive ability test choose lower numbers (Burnham et al., 2009). Mixed evidence was reported in Brañas-Garza et al. (2012): the better performance in the CRT test that measures cognitive reflectiveness is associated with lower numbers in the Guess the number game, whereas the outcome of the Raven test measuring visual reasoning and analytic intelligence surprisingly was not associated with the successful performance in the Guess the number game.
Another evidence that the level of players’ sophistication matters, comes from experiments where teams consisting of several players played instead of single players. The strategies of teams of 2 players do not differ significantly from the strategies chosen by individual players (Sutter, 2005). At the same time, teams consisting of 3 and 4 players outperform individual players (Kocher & Sutter, 2005, Sutter, 2005).
@@
For more experimental and theoretical results on Guess the number games, we refer the reader to one of the surveys (Nagel, 2008, Nagel et al., 2017).
@@
Q5. Can LLM correctly use the analytical solution when choosing their strategy?
@@
Though this question is the most important for us, the nuances that may arise in questions Q1Q6 could affect the interpretation of the results. Therefore, we do not restrict our attention to Q7 solely.
@@
After the initial publication of our working paper (Dagaev et al., 2024), we found several new relevant working papers that we would like to mention. Lu (2024) also runs beauty contest experiments with AI players. Whereas the latter paper focuses on the iterated standard beauty contest game and dynamics of the LLMs decisions, our paper replicates a wide class of one-shot experiments with different groups of competitors. Guo et al. (2024) introduced EconArena environment that allows to organize various competitions between AI models. A beauty contest game is one of the first games that was implemented on EconArena.
@@ 6 Grosskopf & Nagel (2008) 2 2/3 Mean First year undergraduate students majoring in economics
10 Brañas-Garza et al. (2012) 24 1/2 Mean individuals with high CRT score
@@ 3 Results
Replication results for Brañas-Garza et al. (2012) experiment are reported in Table 5. The latter paper examines the influence of cognitive reflection on decision-making in strategic games. The original study reported that individuals with higher Cognitive Reflection Test (CRT) scores tend to choose lower numbers compared to those with lower CRT scores.
@@
Table 5: Replication results for Brañas-Garza et al. (2012) with a LLM player.
@@ 4 Discussion
Q5. Can LLM correctly use the analytical solution when choosing their strategy? (No.)
@@
The most convincing argument against the idea that all LLMs produce almost the same outcome is the behavior of Llama. Llama systematically play much higher numbers than other LLMs (see Tables 26). Moreover, Llama often does not follow the expected comparative statics (see Q3), in contrast to other LLMs.
@@
If one would cast a look at Tables 26, one could notice that the difference between the strategies played by humans and by LLMs is rather large. LLMs tend to overestimate the level of sophistication of their opponents and play lower numbers. The only exception is Llama that systematically plays high numbers that are often higher than the numbers played by humans. However, our results suggest that LLMs (except Llama) are rather successful in identifying the underlying strategic environment and adapting the strategies to the changing set of parameters of the game in the same way as human players do (see Q2 and Q3).
@@ 5 Conclusion
Our results show that 4 out of 5 modern LLMs in consideration (Claude Sonnet, Gemini Flash, GPT-4o, and GPT-4o Mini) systematically behave in a more sophisticated way compared to the participants of the original experiments. The LLMs play lower numbers compared to the numbers played by the human players in a similar setting. The only exception is Llama, it typically plays higher or nearly equal numbers compared to human players’ strategies. Also, Llama exhibited more static strategies, reflecting lower sensitivity to changes in experimental settings. All models failed to play a dominant strategy in a two-player game. These results suggest that while some LLMs are capable of emulating nuanced strategic behavior, their responses are often shaped by their underlying architectures and design priorities.
@@ Acknowledgements
The data supporting the findings of this study is publicly available in Harvard Dataverse: https://doi.org/10.7910/DVN/LURKWJ (accessible upon publication of the manuscript).

### v2

@@ Abstract
A $p$-beauty contest is a wide class of games of guessing the most popular strategy among other players. In particular, guessing a fraction of a mean of numbers chosen by all players is a classic behavioral experiment designed to test iterative reasoning patterns among various groups of people. The previous literature reveals that the level of sophistication of the opponents is an important factor affecting the outcome of the game. Smarter decision makers choose strategies that are closer to theoretical Nash equilibrium and demonstrate faster convergence to equilibrium in iterated contests with information revelation. We replicate a series of classic experiments by running virtual experiments with modern large language models (LLMs) who play against various groups of virtual players. We test how advanced the LLMs’ behavior is compared to the behavior of human players. We show that LLMs typically take into account the opponents’ level of sophistication and adapt by changing the strategy. In various settings, most LLMs (with the exception of Llama) are more sophisticated and play lower numbers compared to human players. Our results suggest that LLMs (except Llama) are rather successful in identifying the underlying strategic environment and adopting the strategies to the changing set of parameters of the game in the same way that human players do. All LLMs still fail to play dominant strategies in a two-player game. Our results contribute to the discussion on the accuracy of modeling human economic agents by artificial intelligence.
Keywords: $p$-beauty contest, guess the number, dominant strategy, AI, LLM
@@ 1 Introduction
The recent appearance of large language models (LLMs) has resulted in numerous attempts to substitute humans with generative agents in various settings (see, for example, Park et al. (2023)). The motivation is simple: in many economic activities, the use of algorithms sooner or later will be cheaper and more productive than the use of labor. Still, at this point, it is not clear to what extent LLM can simulate the human’s behavior. The revelation of differences between the decisions of the participants of economic experiments and LLMs’ decisions in similar circumstances becomes an important challenge. Horton (2023) replicates several classic experiments with LLM players and advocates the use of LLMs as models for ordinary economic agents. Akata et al. (2023) reveal that LLMs underperform in the iterated games that require coordination such as the iterated Battle of Sexes. Brookins & DeBacker (2023) show that LLMs prefer fair decisions in a Dictator game, even more than human players do. Goli & Singh (2024) study the intertemporal preferences of LLMs and figure out that artificial players are less patient than humans. In general, different LLMs have different peculiarities: some are very sensitive to game structure, others have issues with context framing (Lorè & Heydari, 2024).
Our paper contributes to this strand of literature by studying the behavior of LLMs in the classic Guess the number game which belongs to a wider class of $p$-beauty contest games. These $p$-beauty contest games are of particular importance because they arise in various industries where the profit of a firm or the payoff of an individual player depend on the median preferences of all economic agents. The behavior of short-term traders and the resulting asset pricing is regarded as a variant of a beauty contest game (Allen et al., 2006, Cespa & Vives, 2015). On the prediction markets, the desire to guess the most popular opinion is often considered as a behavioral driving force in addition to guessing the fundamental probabilities of an outcome (Marinovic et al., 2011). The beauty contest auction is an important alternative to price-only mechanisms of allocating goods to the bidders (Yoganarasimhan, 2016). Whereas algorithmic trading and betting strategies written by humans have long been used on the markets, the appearance of strategies generated by LLMs is a rather recent phenomenon. Investigation of the LLMs’ performance in a $p$-beauty contest game (both absolute and relative to humans) would help to predict their performance in the above-mentioned markets.
The Guess the Number game tests the ability of a player to make a sequence of conclusions, and the outcome depends on their level of sophistication. The literature provides several important observations about the performance of heterogeneous players in various strategic settings. First, and most obvious, more advanced players reach better outcomes. The latter observation is especially notable in antagonistic pure games of skill. Charness (1981) revealed that the objective quality of moves selected by chess players increases with the player’s skill. Levitt et al. (2011) considered a sequential game Race to 100 which is a pure game of skill. They conclude that the ability to perform backward induction leads to better results in the Race to 100 game. Second (and this is less obvious), the players pay attention to the opponents’ quality. In the Eichberger et al. (2008) experimental study, most of the participants (72%) who play the simultaneous games of strategic complements or strategic substitutes feel that they can predict the actions of a game theorist better than the actions of a granny. The same large majority (72%) prefer to play against a game theorist rather than against a granny. In the Prisoner’s Dilemma game, the high-ability players who learn that their partner is a low-ability one, decrease the level of cooperation (Lambrecht et al., 2024). In antagonistic games, more advanced human players demonstrate greater adaptability to competing environment. For example, in the centipede game, stronger chess players exploit the non-equilibrium play of weaker players (Palacios-Huerta & Volij, 2009). However, Levitt et al. (2011) show that the ability to perform backward induction is not related to ability to stop early in the centipede game and assert that, instead of a centipede game, the pure games of skill should be used to test the role of the level of sophistication. We conclude that in many strategic environments, the skill of the players affects the outcome by allowing them to choose better strategies and/or to show greater adaptability to the opponents. To what extent do LLMs behave like a human? To answer this question, we replicate a series of well-known experiments with human participants playing the Guess the Number game by asking LLMs to play against the same groups of competitors.
@@
The rules of the Guess the number game are as follows. A group of $n$ players simultaneously and independently choose a number between 00 and $100$. Denote by $m$ the mean of all strategies played. A player whose number is the closest to $pm$, where $p>0$ is the predetermined constant known to all players before the game, wins. In case of a tie, all tied players get the corresponding share of the prize. When all real numbers from $[0,100]$ are allowed, for $p\in(0,1)$ there is a unique Nash equilibrium in the model where all players choose 00. In a particular case of $n=2$, choosing 00 is a weakly dominant strategy. If only integer numbers are allowed, there do exist additional Nash equilibria where the players play higher numbers. For example, if $p=0.5$, all strategy profiles where most of the players play $1$ and other players play 00, are additional Nash equilibria. If $p=2/3$, the profile $(1,\ldots,1)$ is the only additional Nash equilibrium.
@@
In a particular case of $n=2$, one could possibly anticipate a significant share of players choosing 00, a weakly dominant strategy. However, this is not the case. Only $10\%$ of undergraduate students and $37\%$ of the audience of economics or psychology decision-making conferences chose 00 (Grosskopf & Nagel, 2008). Also, the mean of the numbers chosen by the professionals (22) is lower than the mean of the numbers chosen by the students (36). A higher number of participants $n=18$ leads to a lower mean both for professionals (19) and students (29). However, in the case of professionals, this difference is not statistically significant (Grosskopf & Nagel, 2008). Rydval et al. (2009) strengthen this finding by identifying that only nearly 1/3 of all participants think in terms of dominant strategies in 2- or 3-person Guess the number-like games with dominant strategies, while 2/3 of all participants fail to identify the strategic properties of the game. In our study, we ask whether LLMs are able to identify the strategic nature of the game.
@@
Cognitive ability is also an important determinant of the outcome of the Guess the number game. Higher cognitive ability may manifest itself through better inductive reasoning, iterative dominance, and level-$k$ thinking (Carpenter et al., 2013). Brocas & Carrillo (2020) designed a variant of the Guess the number game and demonstrated that the equilibrium behavior increases significantly between 5 and 10 years of age (from 17.9% to 61.4%) and stabilizes afterward. Back to the classic variant of the game, players with higher scores in a cognitive ability test choose lower numbers (Burnham et al., 2009) and show faster convergence to equilibrium in iterated experiments (Gill & Prowse, 2016). Mixed evidence was reported in Brañas-Garza et al. (2012): the better performance in the CRT test that measures cognitive reflectiveness is associated with lower numbers in the Guess the number game, whereas the outcome of the Raven test measuring visual reasoning and analytic intelligence surprisingly was not associated with the successful performance in the Guess the number game. Interestingly, high cognitive ability test scorers better respond to the cognitive ability of their opponents (Gill & Prowse, 2016), whereas players whose abilities are below a certain threshold do not adapt strategically to the opponents’ level of sophistication at all Fehr & Huck (2016). The recent Gill et al. (2025) study distinguishes between cognitive ability and judgment. The authors show that whereas high cognitive ability shifts the strategy towards 0, high judgment subjects are less inclined to choose 0, even though their choices on average are lower than low judgment subjects. Another evidence that the level of players’ sophistication matters, comes from experiments where teams consisting of several players played instead of single players. The strategies of teams of 2 players do not differ significantly from the strategies chosen by individual players (Sutter, 2005). At the same time, teams consisting of 3 and 4 players outperform individual players (Kocher & Sutter, 2005, Sutter, 2005). One should also distinguish between cognitive abilities and cognitive effort. Alós-Ferrer & Buckenmaier (2021) propose an experiment where the deliberation time in the Guess the number game serves as a proxy for cognitive effort. The authors find evidence that longer deliberation time is related to playing strategies associated with higher numbers of reasoning steps. After running experiments with two families of games, Georganas et al. (2015) warn us that strategic sophistication varies across different families of games which prevents us from making too strong conclusions and expanding them from the Guess the number game to more general environments.
@@
There are two principal dimensions of the problem in consideration.
  1. 1.
    For a wide range of parameters of a one-shot game, one can compare the strategies of humans and LLMs in a one-shot Guess the number game.
  2. 2.
    For a specific set of parameters, one can compare the strategies of humans and LLMs in the iterated version of the Guess the number game.
While the first approach focuses on the role of changing conditions of the game, the second approach studies learning effects. Our paper follows the former one. We do not observe learning effects neither in humans nor in LLMs. In order to make a fair comparison, we took only the first iterations of the game from the previous literature. Similarly, we never asked LLMs to play the Guess the number game several times in a row during one iteration of an experiment, and we never provided any feedback to LLMs. In contrast, the working paper Lu (2024) follows the second approach.
For more experimental and theoretical results on Guess the number games, we refer the reader to one of the surveys (Nagel, 2008, Nagel et al., 2017). Fan et al. (2024) presents a detailed overview of LLMs research published by 2023.
@@
Q5. Can LLM correctly use the analytical solution when choosing its strategy?
@@
Though this question is the most important for us, the nuances that may arise in questions Q1–Q6 could affect the interpretation of the results. Therefore, we do not restrict our attention to Q7 solely.
@@
Finally, note that we did not organize games between LLMs in this study. We refer those who are interested in tournaments between artificial players to Guo et al. (2024) who introduced EconArena environment that allows to organize various competitions between AI models. A beauty contest game is one of the first games that was implemented on EconArena.
@@ 2 The Experimental Design
For reproducibility purposes we have now detailed all relevant experimental parameters in Table A1 in the Appendix. Specifically, we employed the default settings provided by the respective APIs for each model version—namely, GPT-4o, GPT-4o mini, Gemini-1.5-flash, Claude-3.5-Sonnet, and Llama-3.1-8B-Instruct-Turbo—including default token limits and temperature settings. The data collection process was repeated within a single session environment until 50 valid responses were obtained.
@@ 6 Grosskopf & Nagel (2008) 2 2/3 Mean First year undergraduate students majoring in economics
10 Braùas-Garza et al. (2012) 24 1/2 Mean individuals with high CRT score
@@ 3 Results
Replication results for Braùas-Garza et al. (2012) experiment are reported in Table 5. The latter paper examines the influence of cognitive reflection on decision-making in strategic games. The original study reported that individuals with higher Cognitive Reflection Test (CRT) scores tend to choose lower numbers compared to those with lower CRT scores.
@@
Table 5: Replication results for Braùas-Garza et al. (2012) with a LLM player.
@@ 4 Discussion
Q5. Can LLM correctly use the analytical solution when choosing its strategy? (No.)
@@
The most convincing argument against the idea that all LLMs produce almost the same outcome is the behavior of Llama. Llama systematically play much higher numbers than other LLMs (see Tables 2 –6). Moreover, Llama often does not follow the expected comparative statics (see Q3), in contrast to other LLMs.
@@
If one would cast a look at Tables 2 –6, one could notice that the difference between the strategies played by humans and by LLMs is rather large. LLMs tend to overestimate the level of sophistication of their opponents and play lower numbers. The only exception is Llama that systematically plays high numbers that are often higher than the numbers played by humans. However, our results suggest that LLMs (except Llama) are rather successful in identifying the underlying strategic environment and adapting the strategies to the changing set of parameters of the game in the same way as human players do (see Q2 and Q3).
@@ 5 Conclusion
Our results show that 4 out of 5 modern LLMs in consideration (Claude Sonnet, Gemini Flash, GPT-4o, and GPT-4o Mini) systematically behave in a more sophisticated way compared to the participants of the original experiments. The LLMs play lower numbers compared to the numbers played by the human players in a similar setting. The only exception is Llama, it typically plays higher or nearly equal numbers compared to human players’ strategies. Also, Llama exhibited more static strategies, reflecting lower sensitivity to changes in experimental settings. All models failed to play a dominant strategy in a two-player game. These results suggest that while some LLMs are capable of emulating nuanced strategic behavior, their responses are often shaped by their underlying architectures and design priorities. A limitation of our study is that we did not incorporate an explicit payoff structure in the AI prompts. The original experiments featured varying incentive mechanisms, which may have influenced participants’ choices. Our decision to leave incentives unspecified—while aiming to replicate original conditions—might have affected the observed strategic behavior of the LLMs. Future research should investigate how different incentive specifications influence AI decision-making in strategic settings.
Another limitation is that the alternative to use the same instructions that humans got in the respective experiments, is not feasible. The context matters. For example, participants of a game theory conference understand who they are. In contrast, LLMs need additional information about their competitors and themselves in the prompt. Therefore, we stick to the second-best option of using prompts that are as similar to the original instructions as possible.
@@ Acknowledgements
The data supporting the findings of this study is publicly available in Harvard Dataverse: https://doi.org/10.7910/DVN/LURKWJ.